# Disk patterns, quasi-duality and the uniform bounded diameter conjecture

### v1

@@ 4.2. Extremal length / width for disk patterns
We denote this complementary component of by $\Pi_{\mathcal{P}}$ (see Figure 4.1), and call it the skinning interstice of $\mathcal{P}$.
@@
$\Gamma_{a,b,\mathcal{P}}\coloneqq\{\alpha\colon\alpha\text{ is a proper path in $\Pi_{\mathcal{P}}$ and connects $\partial D_{a}$ to $\partial D_{b}$}\}.$
@@
$\Gamma_{a,b,\mathcal{P}}^{*}\coloneqq\{\alpha\colon\alpha\text{ is proper pathin $\Pi_{\mathcal{P}}$ and separates $\partial D_{a}$ from $\partial D_{b}$}\}.$
@@
$\operatorname{EW}(\Gamma_{a,b,\mathcal{P}})=\frac{1}{\operatorname{EL}(\Gamma_{a,b,\mathcal{P}})}\text{ and }\operatorname{EW}(\Gamma_{a,b,\mathcal{P}}^{*})=\frac{1}{\operatorname{EL}(\Gamma_{a,b,\mathcal{P}}^{*})}.$
@@ 4.3. Circular rectangles
Let $B(0,R)$ and $B(0,R+1)$ be open disks centered at 00 with radius $R$ and $R+1$ respectively. We denote the annulus
@@
Let $D_{i},i=1,2$ denote the two disk components of ^$\widehat{\mathbb{C}}-\mathscr{A}_{R}$. Let $D$ be a disk in $\widehat{\mathbb{C}}$. We say $D$ is admissible in $\mathscr{A}_{R}$ if
@@ 4.4. Circular rectangle approximating $\Pi_{\mathcal{P}}$
Denote the circular arc $\partial\Pi_{\mathcal{P}}\cap\partial D_{a}$ by ¿ $[x_{1},x_{2}]$ Let $\alpha_{i}\subseteq\partial\Pi_{\mathcal{P}}$ so that $\operatorname{Int}(\alpha_{i})$ is the component of $\partial\Pi_{\mathcal{P}}\cap\mathscr{A}_{R}$ connecting $x_{i}$ to $\partial D_{b}$. Let $p(t)$ be the orthogonal projection of $t\in\alpha_{i}$ onto $\partial D_{a}$.
@@
Let $W$ be the circular length of ¿ $[x_{1},x_{2}]$ If $W+10N\leq 2\pi R$, then by the previous claim, $\Pi_{\mathcal{P}}$ is contained in a circular rectangle $\mathscr{R}_{+}$ of circular width $\leq W+10N$. So by Lemma 4.7,
@@ 4.5. Overflow of vertical and horizontal foliations
Let $\alpha$ be an arc in the vertical foliation $\mathcal{F}_{ver,\mathscr{R}_{-}}$. Let $F$ be a hyperbolic face of $\mathcal{R}(P)$. Let $\Pi_{F,\mathcal{P}}$ be the complementary region of ^$\widehat{\mathbb{C}}-\bigcup_{v\in\partial F}D_{v}$ that has empty intersection with the disk pattern $\mathcal{P}$. If $\alpha\cap\Pi_{F,\mathcal{P}}$ connects $\partial D_{v}$ and $\partial D_{w}$, then we add the edge $vw$ to the graph if no such edge exists. Since the arcs in $\mathcal{F}_{ver,\mathscr{R}_{-}}$ do not cross, it is easy to see that the additional edges we add do not cross. Therefore, there are only finitely many edges we can add, and we obtain a graph extension $\mathcal{H}_{ver}$ satisfying the requirement of the lemma. By adding more edges if necessary, we may assume that $\mathcal{H}_{ver}$ is a triangulation of $P$. The construction of $\mathcal{H}_{hor}$ is similar. ∎
@@ 5.1. Lamination on $P$
$\operatorname{EW}(\Gamma_{a,b,\mathcal{P}})>2\text{ and }\operatorname{EW}(\Gamma_{a^{\prime},b^{\prime},\mathcal{P}})>2.$
@@ 5.2. The uniform bound
Let $a,b$ be a pair of non-adjacent vertices of $\partial P$. Suppose that $\{a,b\}\notin\mathcal{L}$. Then $\operatorname{EW}(\Gamma_{a,b,\mathcal{P}})\leq MN^{3}$ for all $\mathcal{P}\in\operatorname{Teich}(\mathcal{G},\omega)$. The lemma follows. ∎

### v2

@@ 4.2. Extremal length / width for disk patterns
We denote this complementary component by $\Pi_{\mathcal{P}}$ (see Figure 4.1), and call it the skinning interstice of $\mathcal{P}$.
@@
[…]
@@
[…]
@@
[…]
@@ 4.3. Circular rectangles
Let $B(0,R)$ and $B(0,R+1)$ be open disks centered at 0 with radius $R$ and $R+1$ respectively. We denote the annulus
@@
Let $D_{i},i=1,2$ denote the two disk components of $\widehat{\mathbb{C}}-\mathscr{A}_{R}$. Let $D$ be a disk in $\widehat{\mathbb{C}}$. We say $D$ is admissible in $\mathscr{A}_{R}$ if
@@ 4.4. Circular rectangle approximating $\Pi_{\mathcal{P}}$
Denote the circular arc $\partial\Pi_{\mathcal{P}}\cap\partial D_{a}$ by ¿ $[x_{1},x_{2}]$ . Let $\alpha_{i}\subseteq\partial\Pi_{\mathcal{P}}$ so that $\operatorname{Int}(\alpha_{i})$ is the component of $\partial\Pi_{\mathcal{P}}\cap\mathscr{A}_{R}$ connecting $x_{i}$ to $\partial D_{b}$. Let $p(t)$ be the orthogonal projection of $t\in\alpha_{i}$ onto $\partial D_{a}$.
@@
Let $W$ be the circular length of ¿ $[x_{1},x_{2}]$ . If $W+10N\leq 2\pi R$, then by the previous claim, $\Pi_{\mathcal{P}}$ is contained in a circular rectangle $\mathscr{R}_{+}$ of circular width $\leq W+10N$. So by Lemma 4.7,
@@ 4.5. Overflow of vertical and horizontal foliations
Let $\alpha$ be an arc in the vertical foliation $\mathcal{F}_{ver,\mathscr{R}_{-}}$. Let $F$ be a hyperbolic face of $\mathcal{R}(P)$. Let $\Pi_{F,\mathcal{P}}$ be the complementary region of $\widehat{\mathbb{C}}-\bigcup_{v\in\partial F}D_{v}$ that has empty intersection with the disk pattern $\mathcal{P}$. If $\alpha\cap\Pi_{F,\mathcal{P}}$ connects $\partial D_{v}$ and $\partial D_{w}$, then we add the edge $vw$ to the graph if no such edge exists. Since the arcs in $\mathcal{F}_{ver,\mathscr{R}_{-}}$ do not cross, it is easy to see that the additional edges we add do not cross. Therefore, there are only finitely many edges we can add, and we obtain a graph extension $\mathcal{H}_{ver}$ satisfying the requirement of the lemma. By adding more edges if necessary, we may assume that $\mathcal{H}_{ver}$ is a triangulation of $P$. The construction of $\mathcal{H}_{hor}$ is similar. ∎
@@ 5.1. Lamination on $P$
[…]
@@ 5.2. The uniform bound
Let $a,b$ be a pair of non-adjacent vertices of $\partial P$. Suppose that $\{a,b\}\notin\mathcal{L}$. Then $\operatorname{EW}(\Gamma_{a,b,\mathcal{P}})\leq MN^{3}$ for all $\mathcal{P}\in\operatorname{Teich}(\mathcal{G},\omega)$. ∎